Please help with this problem?
Solve this absolute value problem.

15 - |2-3x|/2 = 7

Thank you.

Answer by Nate(3500)

- |2 - 3x|/2 = 7
- |2 - 3x| = 14
|2 - 3x| = - 14
No need to further solve because you can not get a negative number from an absolute value.